3x+2y=5 and 2x^2+xy+y^2=14.
Please solve for me I am dying to know how,thanks

Answer by richwmiller(17219)

It is a line intersecting an ellipse (center at (0,0)) in two places
x = -1, y = 4
x = 31/11, y = -19/11